find an angle whose supplement is 2/3 of it?
Answer by stanbon(57377)

find an angle whose supplement is 2/3 of it?
-------------------
Let the angle be "x".
---
Supplement = (2/3)x.
-------------------------
Equation:
x + (2/3)x = 180 degrees
---
(5/3)x = 180
---
x = (3/5)180
---
x = 3*36
---
x = 108 degrees (the angle)
(2/3)x = (2/3)108 = 72 degrees (the other angle)
-----------------------------------------------
Cheers,
Stan H.
===============